draw a plot of the lorentz factor for 0<B<0.9 to see how its value change


1
Expert's answer
2021-06-07T09:35:41-0400

Explanations & Answer


  • "\small \beta" is the ratio between the speed of the object to the speed of light: "\small \beta = \large \frac{v}{c}"
  • And the Lorentz factor is given as "\small \gamma =\large \frac{1}{\sqrt {1-\beta^2}}"
  • Then, its behavior is given as follows,



  • It is visible that the speed of the object approaches the speed of light, Lorentz factor approaches infinity